State TRUE or FALSE justifying your answer with proper reason.



a. 2𝑛^2 + 1 = 𝑂(𝑛^2 )



b. 𝑛^2 (1 + βˆšπ‘›) = 𝑂(𝑛^2 )



c. 𝑛^2 (1 + βˆšπ‘›) = 𝑂(𝑛^2 log 𝑛)



d. 3𝑛^2 + βˆšπ‘› = 𝑂(𝑛 + π‘›βˆšπ‘› + βˆšπ‘›)



e. βˆšπ‘› log 𝑛 = 𝑂(𝑛)

1
Expert's answer
2022-01-24T16:13:54-0500

a.

True

"2𝑛^2 + 1\le 3n^2"


b.

False

"\displaystyle \lim_{n\to \infin} \frac{𝑛^2 (1 + \sqrt𝑛) }{n^2}=\infin"


c.

False

"\displaystyle \lim_{n\to \infin} \frac{𝑛^2 (1 + \sqrt𝑛) }{n^2logn}=\infin"


d.

False

"\displaystyle \lim_{n\to \infin} \frac{ 3𝑛^2 + \sqrt𝑛) }{𝑛 + 𝑛\sqrt𝑛 + \sqrt𝑛}=\infin"


e.

True

"\sqrt𝑛 log 𝑛\le n"